OpenStudy (anonymous):

if A divides B, then the divisors of A also divide B. how ? could somebody help by giving quick explanation plzzz layman's terms please

OpenStudy (mayankdevnani):

this question is a type of transitive relation...

OpenStudy (mayankdevnani):

based on Euclid's lemma.. right.. @sara12345

OpenStudy (anonymous):

hmm idk euclid lemma... i came across this in a solution to puzzle

OpenStudy (anonymous):

consider two numbers 32 and 8 where 8 divides 32 32/8 = 4 Now consider the divisors of 8 which are 2,4 and 8 these also divisors of 32.......... as they all divide 32 Thats it............. Hope this is ok for u

OpenStudy (anonymous):

the six digit number a1989b is ddivisible by 72, determine a and b

OpenStudy (zzr0ck3r):

what it is saying is.. let a = 16 and b = 32 since a devides b, and a = 4*4 or 2*8 or 1*16 we know that 2*4*1*8*18 devide 32

OpenStudy (zzr0ck3r):

wow to weird we almost did the same example

OpenStudy (anonymous):

32 divosors are 16 = 8*2 = 4*2*2 = 2*2*2*2 2, 4, 8 and 16

OpenStudy (anonymous):

of 16 divides 32, then 2, 4, 8 also divide 32 is it

OpenStudy (anonymous):

if *

OpenStudy (zzr0ck3r):

right so since we know 32 devides 64, we know 2,4, 8 devide 64

OpenStudy (anonymous):